# RMAAT: Astrocyte-Inspired Memory Compression and Replay for Efficient Long-Context Transformers

### v1

@@ 1 Introduction
Among these overlooked elements are astrocytes, a type of glial cells increasingly recognized not just for support functions but for their active participation in modulating synaptic transmission, plasticity, and memory processes critical for learning (Gibbs et al., 2008; Bohmbach et al., 2022; Perea et al., 2009; Alberini et al., 2018). Given their established role in modulating temporal information and memory consolidation within biological circuits, we build on the premise that principles derived from astrocyte function are particularly well-suited to addressing the long-range temporal dependency challenges inherent in processing extended sequences. Despite their potential, astrocyte-based computational principles remain severely underexplored in deep learning. This paper introduces the Recurrent Memory Augmented Astromorphic Transformer (RMAAT), an architecture that integrates specific, computationally abstracted astrocyte-inspired mechanisms related to temporal memory processing (inspired by astrocyte long-term effects) and attention modulation (inspired by astrocyte short-term effects) within a recurrent transformer framework. Our goal is to leverage these neuro-glial principles to create an efficient approach for long-context sequence processing. We validate RMAAT’s effectiveness by demonstrating highly competitive results on the full Long Range Arena (LRA) benchmark (Tay et al., 2020) and substantial gains in computational efficiency compared to baselines. The remainder of this paper is organized as follows: Section 2 details our main contributions and positions RMAAT relative to prior work. Section 3 describes the RMAAT model architecture and its bio-inspired components. Section 4 presents experiments and results. Section 5 discusses limitations and concludes the paper.
@@ 2 Related Works and Main Contributions
Significant research addresses the ($O(N^{2})$) complexity and long-context limitations of standard Transformers (Vaswani et al., 2017). Early efficiency improvement efforts focused on sparse or linear attention approximations (e.g., Longformer (Beltagy et al., 2020), Reformer (Kitaev et al., 2020)). Others incorporated recurrence via state caching or compression (e.g., Transformer-XL (Dai et al., 2019), Compressive Transformer (Rae et al., 2019)) and some utilized explicit memory tokens to carry context between segments (e.g., RMT (Bulatov et al., 2022), Memformer (Wu et al., 2020)). More recently, highly efficient architectures like State-Space Models (e.g., S4 (Gu et al., 2021), Mamba (Gu & Dao, 2023)) based on continuous-time systems, and RNN/Transformer hybrids (e.g., RetNet (Sun et al., 2023), RWKV (Peng et al., 2023), GLA (Yang et al., 2023)) employing innovations like retention mechanisms or gating, have achieved strong results through sophisticated architectural and mathematical advancements. However, developing methods that integrate deeper biological principles, particularly for complex functions like long-term memory integration, alongside computational efficiency remains an ongoing challenge. Separately, within biologically-inspired computing, astromorphic approaches have explored leveraging astrocyte principles (Kozachkov et al., 2023), particularly adapting attention mechanisms based on astrocyte nonlinearities and inherent plasticities (Mia et al., 2025). While valuable, these efforts have primarily concentrated on the attention component itself. The potential for utilizing computational principles derived from astrocyte temporal dynamics, such as those involved in long-term plasticity (LTP) related to memory formation and consolidation, to specifically address the challenge of long-range context propagation in sequence models remains largely unexplored. To address these gaps, the main contributions of our work are:
@@
(i) A Distilled Computational Macro Model: We propose and utilize a novel macro model, distilled from detailed computational models of neuron-astrocyte LTP dynamics (Perea et al., 2009; Alberini et al., 2018), which serves as the foundation for RMAAT’s recurrent memory system. (ii) Astrocyte-Inspired Memory Mechanism: We introduce an integrated memory system featuring persistent memory tokens inspired by astrocyte processes. This system incorporates a novel memory retention factor, derived from the macro model, to achieve biologically-motivated context compression and propagation, differing significantly from architectures reliant on externally managed memory (Bulatov et al., 2022; Wu et al., 2020). (iii) Efficient AMRB Training Algorithm: We propose the Astrocytic Memory Replay Backpropagation (AMRB) algorithm, enabled by the model’s memory structure, which significantly reduces the memory footprint and computational overhead compared to standard BPTT or chunk-based backpropagation for recurrent training.
@@ 3.1 Foundational Computational Neuroscience Model
The dynamics governed by Equation 3, representing the integration of synaptic history ($s_{ij}$) over longer timescales via the $p^{l}_{ij}$ variable, provides the conceptual foundation for our subsequent developments. Specifically, we distill the principles captured by these LTP dynamics into a computational Macro Model (Contribution 1), which then informs the design of our Astrocyte-Inspired Memory Mechanism (Contribution 2), including the derivation of a memory retention factor for context compression of memory tokens. The detailed derivation, simulation results showing the characteristic behavior, and implementation of this memory system are presented in Section 3.3.
These equations represent the core conceptual model adapted from foundational works (Kozachkov et al., 2025; Wade et al., 2011). Full model details, parameters, and the underlying neural dynamics governing $x_{i},x_{j}$ are provided in Appendix A. The subsequent sections detail how RMAAT translates these principles into a computational architecture.
@@ 3.2.1 Segmented Processing and Biologically Inspired Memory Tokens
To address the quadratic complexity bottleneck of standard self-attention over long sequences, RMAAT adopts a segmented processing approach. The input sequence is divided into non-overlapping, contiguous segments of a manageable maximum length $N_{seg}$. The core RMAAT layers process these segments sequentially, rather than operating on the entire sequence at once. A key element enabling long-range dependency modeling across these segments is the incorporation of dedicated Memory Tokens. Inspired by the capacity of biological systems, particularly astrocyte networks, to maintain and integrate information over extended periods (as abstracted in Sec 3.1), these memory tokens serve as a persistent, evolving state. Let the set of $M$ memory tokens at the start of processing segment $t$ be denoted by $mem_{t}$. These tokens are processed alongside the actual input tokens $x_{t}$ within the segment using the mechanisms described below (Sec 3.2.2 and Sec 3.2.3). The output representations corresponding to these memory tokens after processing segment $t$ form the updated memory state, $mem_{t+1}$, which is then passed as the input memory to segment $t+1$. This recurrent flow, conceptually illustrated in Figure 1, allows contextual information within the memory tokens to propagate across segments. This mechanism differs from approaches like RMT (Bulatov et al., 2022) or Memformer (Wu et al., 2020), which often rely on externally managed memory mechanisms or specific architectural additions for memory updates. In RMAAT, the update of these memory tokens are intrinsically linked to the bio-inspired dynamics derived from our computational macro model (detailed in Sec 3.3, involving a dynamically derived retention factor), aiming for a more integrated and computationally distinct approach to memory management. The processing within each segment, which updates both sequence and memory token representations, relies on the astromorphic attention mechanism described below.
Figure 2: Overview of the Astromorphic Transformer architecture. This diagram illustrates the integration of bioplausible bidirectional feedback mechanisms within a two-layered neuron-astrocyte network, emulating the Self-Attention of the transformer encoder. The synaptic weights $W_{K}$, $W_{Q}$, and $W_{V}$ are corresponding to key ($K$), query ($Q$) and value ($V$) of the transformer which are activated once the input ($X$) is received. Write Mode (marked in brown): Hebbian plasticity ($H$) between the pre and postneurons is denoted by $H_{neuron}$ and that between the postneuron and astrocyte is denoted by $H_{astro}$ (dashed lines represent the bidirectional connections among the elements of the tripartite synapse). $H_{astro}$ calculation involves the astrocytic parameter, $R$ which is computed based on the spatial positions of neurons quantified by $T_{ijkl}$. The calcium response of the astrocyte ($C$) is encoded from the $Ca^{2+}$ concentration in the astrocyte evoked by the presynaptic action potential ($K$). Read Mode (marked in red): Presynaptic plasticity ($P$) is decoded by the presynaptic action potential ($Q$) and the final weight $H\odot P$ defines the synaptic weight between the hidden ($h$) and the output layer ($L$).
@@ 3.2.2 Astromorphic Attention Mechanism
The Write Mode then computes effective synaptic weights and states within this network structure, encoding context based on Hebbian principles and astrocyte modulation. These represent learned parameters or aggregated states within the network. Specifically: The Neuronal Hebbian Weight component ($H_{neuron}\in\mathbb{R}^{m\times d}$) captures the direct correlation between activated keys $\phi(K)$ (hidden layer activations) and values $V$ (output layer), representing baseline Hebbian plasticity summed across the $N$ tokens (conceptually linked to the $\theta(x_{i})\theta(x_{j})$ co-activation term in Eq. 1). This models the connection strength between the hidden (presynaptic) and output (postsynaptic) layers based on direct neuron-neuron interaction. The Astrocyte-Modulated Hebbian Weight component ($H_{astro}\in\mathbb{R}^{m\times d}$) incorporates the astrocyte’s modulatory influence, specifically integrating relative positional information (conceptually linked to the astrocyte modulation term $\psi(p_{ij}^{s})$ in Eq. 1, where $p_{ij}^{s}$ dynamics are spatially modulated). Building on prior astromorphic transformer work (Mia et al., 2025), it uses the activation of a relative positional encoding matrix $R$ (astrocytic parameter), $\phi(R)$, derived from STP dynamics (detailed in Sec 3.2.3)to represent the influence of relative positioning. This models how astrocytes modulate the hidden-to-output layer connection based on spatial context. Concurrently, the Presynaptic State ($g\in\mathbb{R}^{1\times m}$) abstracts the non-linear astrocyte response (e.g., calcium dynamics, denoted by $C\sim Ca^{2+}$ in Figure 2) to the cumulative presynaptic (key) activity $\phi(K)$ within the segment. It aggregates the activated keys $\phi(k_{t})$ (where $\phi(k_{t})$ is the $t$-th row of $\phi(K)$) over the segment length $N$ and applies a non-linearity controlled by parameter $\alpha$. This state $g$ encodes recent activation history in the hidden layer, influenced by astrocyte dynamics ($\alpha$), relevant for feedback modulation in the Read Mode. These three components, representing learned weights ($H_{neuron},H_{astro}$) and an aggregated state ($g$) within the network, are calculated as:
@@ 3.2.3 Biological Grounding of Relative Positional Encoding by STP Dynamics
Our computational neuroscience simulations (Sec 3.1, Appendix C) investigate spatial interactions among astrocyte processes. These simulations incorporate the distance-dependent coupling tensor $T_{ijkl}$ (Eq. 2), reflecting how influence between processes diminishes with distance, akin to biological signaling like calcium diffusion (Wade et al., 2011; De Pittà et al., 2009). A key outcome is that this coupling naturally induces spatially modulated activity patterns in astrocyte processes ($p^{s}_{ij}$): those near activity centers exhibit higher peak and sustained activity due to stronger integrated neighborly coupling compared to peripheral processes. This inherent encoding of spatial relationships within simulated STP dynamics provides a strong biological rationale for incorporating a similar distance-based relative positional information scheme in our Astromorphic Attention. We translate this observed principle into the Astrocyte-Modulated Hebbian Weight ($H_{astro}$) component via the term $\phi(R)$, using the positional encoding matrix $R$ (computed as described in the first paragraph). Calculating $H_{astro}=\frac{1}{m}\phi(R)^{T}V$ (Eq. 4) thus integrates a form of spatial context whose use is directly motivated by its analogy to simulated astrocyte STP behavior. This offers a flexible, learnable, and biologically-grounded method for incorporating relative positional context, distinct from standard approaches lacking this neuro-glial justification. Having addressed this spatially-informed component of attention, we now turn to the temporal memory mechanisms essential for processing long sequences.
@@ 3.3 Astrocyte-Inspired Memory Mechanism
Computational Macro Model (Contribution $1$): We leverage the detailed computational neuroscience model of neuron-astrocyte interactions (Sec 3.1) to understand the principles underlying LTP. Simulations of this detailed model reveal essential characteristics of the LTP-related state ($p^{l}_{ij}$): gradual integration of information over successive Short-Term Plasticity (STP) cycles, continuous accumulation across these cycles, and eventual saturation. Figure 3 illustrates this simulated behavior for a $3\times 3$ neuron network over $300$ seconds (encompassing six $50$s STP cycles). We distill these observed characteristics—temporal integration and saturation—into a guiding computational macro model principle that forms the foundation for RMAAT’s persistent memory mechanism.
Memory Retention Factor (Contribution $2$): The saturation dynamic inherent in the macro model principle—reflecting finite biological resources like intracellular calcium concentration—inspires a mechanism for context compression. We derive a novel Memory Retention Factor from this principle by analyzing the accumulation rate within the simulated LTP curve (like the one shown for $6$ cycles in Figure 3).
The determination of the factor begins by normalizing the total information capacity—representing the maximum integrated value the LTP state reaches at saturation (conceptually, the area under the dashed curve in Figure 3 until it flattens)—to $1$ unit. We want to find out what fraction of this total capacity is contributed or updated by each segment (STP cycle). To determine this relationship generally, we can run simulations of the computational model assuming various anticipated total sequence lengths (i.e., total number of segments). For any given assumed total length (e.g., $T$ segments), we simulate the LTP dynamics and measure the increase in the LTP state during each segment t=1,2,,Tt=1,2,...,T. We then express the increase during a specific segment $t$ as a fraction of the total normalized capacity ($1$ unit). This fraction yields the Memory Retention Factor for segment $t$ when the total anticipated sequence length is $T$. This process allows us to determine the retention factors for sequences of arbitrary anticipated lengths.
@@
Simulations indicate that each subsequent segment contributes a diminishing fraction to the total memory capacity, with this compressive effect intensifying for sequences anticipated to be longer. Figure 4 illustrates this principle by showing the calculated factors for example total sequence lengths of $2$, 44, 66, and $8$ segments. This derived factor enables adaptive, bio-inspired compression: the model adjusts how much it compresses information based on the expected length of the sequence, mimicking biological resource limits, which leads to older information being gradually compressed more to accommodate newer information.
Application to Memory Tokens: This Memory Retention Factor is applied directly to RMAAT’s persistent Memory Tokens $mem_{t}\in\mathbb{R}^{M\times d}$ (Sec 3.2.1), which carry context across segments. As tokens are updated within a segment via Astromorphic Attention to produce the next state $mem_{t+1}$, the factor corresponding to the current segment number and total sequence length (Figure 4) scales the updated state (e.g., $mem_{t+1}=\text{RetentionFactor}(t,\text{TotalSegments})\times mem_{t+1}^{\prime}$; See Appendix D for details). This implements the adaptive compression dictated by the LTP macro model principle, ensuring memory remains bounded by gradually compressing older information. This contrasts with architectures like RMT (Bulatov et al., 2022) that often rely on fixed-size external memory slots updated via standard mechanisms, lacking this specific bio-inspired, adaptive compression rationale derived from LTP dynamics. This integrated, astrocyte-inspired memory system not only manages long-range context efficiently but also enables the resource-efficient AMRB training algorithm detailed next (Section 3.4).

### v2

# arXiv:2601.00426v2[cs.NE] 28 Feb 2026

@@ 1 Introduction
Among these overlooked elements are astrocytes, a type of glial cell increasingly recognized not just for support functions but for their active participation in modulating synaptic transmission, plasticity, and memory processes critical for learning (Gibbs et al., 2008; Bohmbach et al., 2022; Perea et al., 2009; Alberini et al., 2018). Given their established role in modulating temporal information and memory consolidation within biological circuits, we build on the premise that principles derived from astrocyte function are particularly well-suited to addressing the long-range temporal dependency challenges inherent in processing extended sequences. Despite their potential, astrocyte-based computational principles remain severely underexplored in deep learning.
This paper introduces the Recurrent Memory Augmented Astromorphic Transformer (RMAAT), an architecture that integrates specific, computationally abstracted astrocyte-inspired mechanisms related to temporal memory processing (inspired by astrocyte long-term effects) and attention modulation (inspired by astrocyte short-term effects) within a recurrent transformer framework. Our goal is to leverage these neuro-glial principles to create an efficient approach for long-context sequence processing. Within this emerging line of work, foundational studies have shown that tripartite synapses can implement Transformer self-attention, validated via weights extracted from pre-trained networks (Kozachkov et al., 2023), and have developed theoretical models of neuron–astrocyte associative memory and capacity scaling (Kozachkov et al., 2025). Subsequent Astromorphic Transformer architectures (Mia et al., 2025) have shown that these principles can be instantiated in standard-scale machine learning models. Building on this trajectory, our work scales these astrocyte-inspired mechanisms to a demanding long-context benchmark (LRA), where they yield competitive performance and improved memory efficiency in the long-sequence regime. The remainder of this paper is organized as follows: Section 2 details our main contributions and positions RMAAT relative to prior work. Section 3 describes the RMAAT model architecture and its bio-inspired components. Section 4 presents experiments and results. Section 5 discusses limitations and concludes the paper.
@@ 2 Related Works and Main Contributions
Significant research addresses the ($O(N^{2})$) complexity and long-context limitations of standard Transformers (Vaswani et al., 2017). Early efficiency improvement efforts focused on sparse or linear attention approximations (e.g., Longformer (Beltagy et al., 2020), Reformer (Kitaev et al., 2020)). Others incorporated recurrence via state caching or compression (e.g., Transformer-XL (Dai et al., 2019), Compressive Transformer (Rae et al., 2019)) and some utilized explicit memory tokens to carry context between segments (e.g., RMT (Bulatov et al., 2022), Memformer (Wu et al., 2020)). More recently, highly efficient architectures like State-Space Models (e.g., S4 (Gu et al., 2021), Mamba (Gu & Dao, 2023)) based on continuous-time systems, and RNN/Transformer hybrids (e.g., RetNet (Sun et al., 2023), RWKV (Peng et al., 2023), GLA (Yang et al., 2023)) employing innovations like retention mechanisms or gating, have achieved strong results through sophisticated architectural and mathematical advancements. However, developing methods that integrate deeper biological principles, particularly for complex functions like long-term memory integration, alongside computational efficiency remains an ongoing challenge. Separately, within biologically-inspired computing, astromorphic approaches have explored leveraging astrocyte principles (Kozachkov et al., 2023), particularly adapting attention mechanisms based on astrocyte nonlinearities and inherent plasticities (Mia et al., 2025). While valuable, these efforts have primarily concentrated on the attention component itself. The potential for utilizing computational principles derived from astrocyte temporal dynamics, such as those involved in long-term plasticity (LTP) related to memory formation and consolidation, to specifically address the challenge of long-range context propagation in sequence models remains largely unexplored. Complementary lines of work have investigated neuromodulated Hebbian plasticity in RNNs (Miconi et al., 2020; Duan et al., 2023); in contrast, our model fixes the qualitative modulation structure and timescales from a neuron–astrocyte network simulation and only learns the downstream architectural parameters. To address the above-mentioned gaps, the main contributions of our work are:
@@
(i) A Distilled Computational Macro Model: We propose and utilize a novel macro model, distilled from detailed computational models of neuron-astrocyte LTP dynamics (Perea et al., 2009; Alberini et al., 2018), which serves as the foundation for RMAAT’s recurrent memory system. (ii) Memory Retention Factor for Segment-Based Processing: Building on the macro model (i), we derive a novel Memory Retention Factor that bridges the biological abstraction to RMAAT’s recurrent architecture, instantiating the macro model’s saturation dynamics as a concrete compression schedule for segment-based processing of memory tokens. This factor achieves biologically-motivated context compression, differing significantly from architectures reliant on externally managed memory (Bulatov et al., 2022; Wu et al., 2020). (iii) Efficient AMRB Training Algorithm: We propose the Astrocytic Memory Replay Backpropagation (AMRB) algorithm, enabled by the model’s memory structure, which significantly reduces the memory footprint and computational overhead compared to standard BPTT or chunk-based backpropagation for recurrent training.
@@ 3.1 Foundational Computational Neuroscience Model
The dynamics governed by Equation 3, representing the integration of synaptic history ($s_{ij}$) over longer timescales via the $p^{l}_{ij}$ variable, provides the conceptual foundation for our subsequent developments. Specifically, we distill the principles captured by these LTP dynamics into a computational Macro Model (Contribution 1), which we then operationalize for segment-based sequence processing by deriving a Memory Retention Factor (Contribution 2) that instantiates the macro model’s saturation curve as a concrete compression schedule for RMAAT’s recurrent memory tokens. The detailed derivation, simulation results showing the characteristic behavior, and implementation of this memory system are presented in Section 3.3.
The subsequent sections detail how RMAAT translates these principles into a computational architecture, moving from biophysical simulation to a macro-model of STP/LTP and then to concrete mechanisms for within-segment attention modulation (STP-like) and cross-segment recurrent memory integration (LTP-like) (Debanne & Inglebert, 2023; Bicknell & Latham, 2024; Stasenko & Kazantsev, 2023).
@@ 3.2.1 Segmented Processing and Biologically Inspired Memory Tokens
To address the quadratic complexity bottleneck of standard self-attention over long sequences, RMAAT adopts a segmented processing approach. The input sequence is divided into non-overlapping, contiguous segments of a manageable maximum length $N_{seg}$. The core RMAAT layers process these segments sequentially, rather than operating on the entire sequence at once. A key element enabling long-range dependency modeling across these segments is the incorporation of dedicated Memory Tokens. Inspired by the capacity of biological systems, particularly astrocyte networks, to maintain and integrate information over extended periods (as abstracted in Sec 3.1), these memory tokens serve as a persistent, evolving state. Let the set of $M$ memory tokens at the start of processing segment $t$ be denoted by $mem_{t}$. These tokens are processed alongside the actual input tokens $x_{t}$ within the segment using the mechanisms described below (Sec 3.2.2 and Sec 3.2.3). The output representations corresponding to these memory tokens after processing segment $t$ form the updated memory state, $mem_{t+1}$, which is then passed as the input memory to segment $t+1$. This recurrent flow, conceptually illustrated in Figure 1, allows contextual information within the memory tokens to propagate across segments. This mechanism differs from approaches like RMT (Bulatov et al., 2022) or Memformer (Wu et al., 2020), which often rely on externally managed memory mechanisms or specific architectural additions for memory updates. In RMAAT, the updates to these memory tokens are intrinsically linked to the bio-inspired dynamics derived from our computational macro model (detailed in Sec 3.3, involving a dynamically derived retention factor), aiming for a more integrated and computationally distinct approach to memory management. The processing within each segment, which updates both sequence and memory token representations, relies on the astromorphic attention mechanism described below.
@@ 3.2.2 Astromorphic Attention Mechanism
The Write Mode then computes effective synaptic weights and states within this network structure, encoding context based on Hebbian principles and astrocyte modulation. These represent learned parameters or aggregated states within the network. Specifically: The Neuronal Hebbian Weight component ($H_{neuron}\in\mathbb{R}^{m\times d}$) captures the direct correlation between activated keys $\phi(K)$ (hidden layer activations) and values $V$ (output layer), representing baseline Hebbian plasticity summed across the $N$ tokens (conceptually linked to the $\theta(x_{i})\theta(x_{j})$ co-activation term in Eq. 1).
Figure 2: Overview of the Astromorphic Attention mechanism, detailing the flow from input to output. This architecture replaces standard self-attention with a bio-inspired, linear-complexity mechanism that emulates the function of a tripartite synapse. The process unfolds in three main steps. (1) Input Projection: The input sequence $X$, which includes both the segment’s tokens and the recurrent memory tokens, is linearly projected into Key ($K$), Query ($Q$), and Value ($V$) matrices using learnable weight matrices $W_{K}$, $W_{Q}$, and $W_{V}$. (2) Write Mode (Context Encoding, brown path): This phase aggregates and encodes contextual information across the entire input segment, modeling how astrocytes integrate signals over a local neighborhood. It computes two Hebbian weight matrices: the Neuronal Hebbian Weight ($H_{neuron}$), capturing direct key-value correlations, and the Astrocyte-Modulated Hebbian Weight ($H_{astro}$), which is modulated by the astrocytic parameter RR—a learnable matrix that abstracts the principles of distance-dependent spatial coupling quantified by the tensor $T_{ijkl}$ in the underlying neuroscience model (Sec. 3.2.3)—and introduces spatial context. Together, these Hebbian weights define effective synaptic connections from the hidden layer ($h$) to the output layer ($L$) depicted in the figure. Concurrently, the mechanism computes a presynaptic state $g$, which represents the astrocyte’s integrated response to overall neuronal activity and computationally abstracts the dynamics of intracellular calcium concentration ($Ca^{2+}$). (3) Read Mode (Context Retrieval, red path): This phase uses the encoded context to generate the final output for each token, modeling astrocyte-mediated feedback. First, each token’s unique Query vector is used to compute a dynamic Presynaptic Plasticity Feedback Factor ($P$), which acts as a query-specific feedback signal. This factor then modulates the combined Hebbian weights ($H=H_{neuron}+H_{astro}$), and the resulting modulated weight is applied to the queries to produce the final output $L$ at the output layer. The full equations are detailed in Section 3.2.2.
This models the connection strength between the hidden (presynaptic) and output (postsynaptic) layers based on direct neuron-neuron interaction. The Astrocyte-Modulated Hebbian Weight component ($H_{astro}\in\mathbb{R}^{m\times d}$) incorporates the astrocyte’s modulatory influence, specifically integrating relative positional information (conceptually linked to the astrocyte modulation term $\psi(p_{ij}^{s})$ in Eq. 1, where $p_{ij}^{s}$ dynamics are spatially modulated). Building on prior astromorphic transformer work (Mia et al., 2025), it uses the activation of a relative positional encoding matrix $R$ (astrocytic parameter), $\phi(R)$, derived from STP dynamics (detailed in Sec 3.2.3) to represent the influence of relative positioning. This models how astrocytes modulate the hidden-to-output layer connection based on spatial context. Concurrently, the Presynaptic State ($g\in\mathbb{R}^{1\times m}$) abstracts the non-linear astrocyte response (e.g., calcium dynamics, denoted by $C\sim Ca^{2+}$ in Figure 2) to the cumulative presynaptic (key) activity $\phi(K)$ within the segment. It aggregates the activated keys $\phi(k_{t})$ (where $\phi(k_{t})$ is the $t$-th row of $\phi(K)$) over the segment length $N$ and applies a non-linearity controlled by parameter $\alpha$. This state $g$ encodes recent activation history in the hidden layer, influenced by astrocyte dynamics ($\alpha$), relevant for feedback modulation in the Read Mode. These three components, representing learned weights ($H_{neuron},H_{astro}$) and an aggregated state ($g$) within the network, are calculated as:
@@ 3.2.3 Biological Grounding of Relative Positional Encoding by STP Dynamics
Our computational neuroscience simulations (Sec 3.1, Appendix C) investigate spatial interactions among astrocyte processes. These simulations incorporate the distance-dependent coupling tensor $T_{ijkl}$ (Eq. 2), reflecting how influence between processes diminishes with distance, akin to biological signaling like calcium diffusion (Wade et al., 2011; De Pittà et al., 2009) (the $T_{ijkl}$-mediated coupling effectively defines “many-neuron synapses” that couple activity across multiple synapses within an astrocyte’s domain Kozachkov et al. (2025)). In our simulations, this same coupling produces spatially structured activity in the astrocyte processes ($p^{s}_{ij}$): processes near activity centers exhibit higher peak and more sustained activity than peripheral ones, reflecting their stronger integrated neighborly coupling.
This inherent encoding of spatial relationships within simulated STP dynamics provides a strong biological rationale for incorporating a similar distance-based relative positional information scheme in our Astromorphic Attention. We translate this observed principle into the Astrocyte-Modulated Hebbian Weight ($H_{astro}$) component via the term $\phi(R)$, using the positional encoding matrix $R$ (computed as described in the first paragraph). Calculating $H_{astro}=\frac{1}{m}\phi(R)^{T}V$ (Eq. 4) thus integrates a form of spatial context whose use is directly motivated by its analogy to simulated astrocyte STP behavior. This offers a flexible, learnable, and biologically-grounded method for incorporating relative positional context, distinct from standard approaches lacking this neuro-glial justification. Having addressed this spatially-informed component of attention, we now turn to the temporal memory mechanisms essential for processing long sequences.
@@ 3.3 Astrocyte-Inspired Memory Mechanism
Computational Macro Model (Contribution $1$): We leverage the detailed computational neuroscience model of neuron-astrocyte interactions (Sec 3.1) to understand the principles underlying LTP. Simulations of this detailed model reveal essential characteristics of the LTP-related state ($p^{l}_{ij}$): gradual integration of information over successive Short-Term Plasticity (STP) cycles, continuous accumulation across these cycles, and eventual saturation. Figure 3 illustrates this simulated behavior for a $3\times 3$ neuron network over $300$ seconds (encompassing six $50$s STP cycles). We distill these observed characteristics—temporal integration and saturation—into a computational macro model that captures the emergent dynamics of LTP-based memory consolidation at an abstract level, independent of specific architectural choices.
From Macro Model to ML Architecture (Contribution $2$): While the macro model captures the biological principle of saturating memory integration, it does not directly prescribe how to implement this within a recurrent transformer processing sequences segment-by-segment. Contribution 2 bridges this gap by deriving a Memory Retention Factor that instantiates the macro model’s saturation curve as a concrete compression schedule for RMAAT’s memory tokens in segment-based processing.
@@
To construct this factor, we normalize the macro model’s total memory capacity (the integrated LTP signal at saturation) to $1$ unit, then compute each segment’s fractional contribution. Formally, for a sequence with $T$ segments, the Memory Retention Factor for segment $t$ is:
$\text{RetentionFactor}(t,T)=\frac{\Delta p^{l}_{t}}{\sum_{i=1}^{T}\Delta p^{l}_{i}}$ (6)
where $\Delta p^{l}_{t}$ represents the incremental increase in the simulated LTP state during segment $t$, obtained by running the LTP macro model (Eq. 3) for $T$ segments and measuring $\Delta p^{l}_{t}=p^{l}(t\cdot\tau_{cycle})-p^{l}((t-1)\cdot\tau_{cycle})$, with $\tau_{cycle}$ being the duration of one STP cycle. Figure 4 illustrates the resulting factors for sequence lengths of $2$, 44, 66, and $8$ segments, showing that each subsequent segment contributes a diminishing fraction—enabling adaptive compression where the model adjusts retention based on anticipated sequence length, mimicking biological resource constraints.
Application to Memory Tokens: This Memory Retention Factor is applied directly to RMAAT’s persistent Memory Tokens $mem_{t}\in\mathbb{R}^{M\times d}$ (Sec 3.2.1), which carry context across segments. As tokens are updated within a segment via Astromorphic Attention to produce the next state $mem_{t+1}$, the factor corresponding to the current segment number and total sequence length (Figure 4) scales the updated state (e.g., $mem_{t+1}=\text{RetentionFactor}(t,\text{TotalSegments})\times mem_{t+1}^{\prime}$; see Algorithm 1 in Section 3.4 for the full pseudocode and Appendix D for additional implementation notes). This implements the adaptive compression dictated by the LTP macro model principle, ensuring memory remains bounded by gradually compressing older information. This contrasts with architectures like RMT (Bulatov et al., 2022) that often rely on fixed-size external memory slots updated via standard mechanisms, lacking this specific bio-inspired, adaptive compression rationale derived from LTP dynamics. This integrated, astrocyte-inspired memory system not only manages long-range context efficiently but also enables the resource-efficient AMRB training algorithm detailed next (Section 3.4).